OpenStudy (anonymous):

write the equation in standard form x^2+y^2+8x-2y+13=0

OpenStudy (amistre64):

completeing a few squareaa

OpenStudy (amistre64):

+16, +1 (x+4)^2 + (y-1)^2 +30 = 0

OpenStudy (amistre64):

opps...-16-1` `

OpenStudy (amistre64):

13-16-1 = -4

OpenStudy (anonymous):

thank U

OpenStudy (amistre64):

x+4)^2 + (y-1)^2 = 4 youre welcome